OpenStudy (anonymous):

solve this equation show all work 2x-4=15x+9

OpenStudy (gabbyalicorn):

Hi there! Do you know what step to do first? If not, no worries! ^u^

OpenStudy (anonymous):

no i have no idea where to start

OpenStudy (gabbyalicorn):

Okay, that's fine. :) We first subtract 15x from both sides.

OpenStudy (gabbyalicorn):

2x−4−15x=15x+9−15x ^ ^ −13x−4=9

OpenStudy (anonymous):

so next do you minus 4 from both sides

OpenStudy (gabbyalicorn):

You are very close! You actually add 4 to both sides.

OpenStudy (gabbyalicorn):

−13x−4+4=9+4 ^ ^ −13x=13

OpenStudy (anonymous):

then you divide?

OpenStudy (gabbyalicorn):

Yes! :D Do you know what you divide by?

OpenStudy (anonymous):

13?

OpenStudy (gabbyalicorn):

So close! It's actually -13 −13x/−13=13−/13 which equals....

OpenStudy (gabbyalicorn):

sorry, i though I posted that xD ^

OpenStudy (anonymous):

-1 would be the answer

OpenStudy (gabbyalicorn):

Yes, yes, yes!

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Can someone help me? A motorboat travels 332 in 4 hours going upstream 555 and in 5 hours going downstream. What is the speed of the boat in still water, and what is the speed of the current?

OpenStudy (gabbyalicorn):

np